# Generative model for financial time series trained with MMD using a signature kernel

### v2

@@ Abstract
Generating synthetic financial time series data that accurately reflects real-world market dynamics holds tremendous potential for various applications, including portfolio optimisation, risk management, and large scale machine learning. We present an approach for training generative models for financial time series using the maximum mean discrepancy (MMD) with a signature kernel. Our method leverages the expressive power of the signature transform to capture the complex dependencies and temporal structures inherent in financial data. We employ a moving average model to model the variance of the noise input, enhancing the model’s ability to reproduce stylised facts such as volatility clustering. Through empirical experiments on S&P 500 index data, we demonstrate that our model effectively captures key characteristics of financial time series and outperforms a comparable GAN-based approach. In addition, we explore the application of the synthetic data generated to train a reinforcement learning agent for portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
@@ 1 Introduction
The availability of high-quality, realistic synthetic data for financial time series holds immense potential for various quantitative finance applications. However, capturing the complex and nuanced characteristics of financial markets, such as volatility clustering, fat tails, and long-range dependencies, poses a significant challenge for traditional generative models. Recent advances in deep learning have introduced promising techniques, including Variational Autoencoders (VAEs) introduced in [18] and Generative Adversarial Networks (GANs) pioneered in [12], for generating synthetic financial data.
This paper proposes a generative model based on the maximum mean discrepancy (MMD) as defined in [13] with the use of signature kernels proposed by [19] to address these challenges. MMD provides a framework for comparing probability distributions, enabling effective training of the generative model. The signature kernel, derived from the signature transform, offers a powerful tool for capturing nuances of financial time series data. We further enhance the model’s ability to represent stylised facts by employing a moving average model to generate structured noise input, leading to more realistic synthetic data generation.
To demonstrate the effectiveness of our approach, we compare the characteristics of the generated sequence using our MMD approach against an alternative that uses causal optimal transport (COT), which is a type of Wasserstein distance that respects causality (see [1]). We find that although both models are generally able to capture stylised facts of financial time series, there are some improvements in the statistical aspects of the MMD generated sequence. Furthermore, we explore the application of the generated synthetic data for training a reinforcement learning agent on the task of portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
The sections are organised as follows. Section 2 provides an overview of related works in the field of generative models for financial time series that use neural networks. Section 3 details the technical underpinnings behind the methodology. We describe our proposed generative model and the modelling of the noise variance in Section 4. Section 5 presents the experimental results, comparing our approach with a GAN-based model trained using COT and evaluating the performance of the synthetic data generated for portfolio management using reinforcement learning. We also introduce a method to add robustness to the generative model in this section. Finally, Section 6 concludes the paper and discusses potential future research directions.
@@ 2 Related Works
Generative models are a type of statistical model that can be used to generate new data instances. They are especially useful in the financial domain, where they can be used to generate financial time series data. Financial time series data is a sequence of values that are measured at regular intervals, such as daily, weekly, or monthly. Generative models can be used to generate new time series data that is similar to the original data. This is useful for a variety of applications, including portfolio optimisation, risk management, and privacy protection. In addition, generative models can be used to backtest trading strategies or serve as training data for machine learning models. Traditional statistical methods for modelling financial time series include autoregressive models, GARCH models, and stochastic volatility models. We focus instead on related works that also use neural networks as the backbone for the generative model.
@@
One approach is the use of Variational Autoencoders (VAEs) where the generative model is trained to maximise the evidence lower bound (ELBO) of the reference data as introduced in [18]. In [3], the authors trained VAEs to reproduce the log signatures of the reference data with a diagonal Gaussian distribution for the latent space. The choice of log signature is motivated by the fact that it is a unique feature transformation that captures the key characteristics of the path. The challenge lies in inverting the log signature to the original path, which is highly nontrivial (see e.g. [27]).
@@
An even more popular approach is the framework of Generative Adversarial Networks (GANs) pioneered in [12] to train neural networks generators of financial time series. An example is the QuantGAN from [37], which focuses on the choice of neural network architecture. In particular, an architecture called Temporal Convolutional Networks (TCN) was used to model drift and volatility processes in order to generate samples. One common issue with the GAN framework is that it is known to be exhibit instabilities during training. The authors of QuantGAN report that "training was very irregular and did not converge to a local optimum of the objective function", hence the need for model selection across multiple parameters saved during training. The QuantGAN was able to capture the stylised facts of financial time series such as fat tails and volatility clustering.
@@
There are other GAN type approaches such as the COT-GAN from [40], which uses a distance metric, calculated with Causal Optimal Transport (COT) or adapted Wasserstein (see [1]), as the discriminator. This falls under the GAN framework as the COT distance is calculated using an optimisation based formulation that uses neural networks as function approximators. Although this particular work was not applied to financial time series, it is an example of a GAN based approach that uses a distance metric as the discriminator exhibiting better stability during training.
Our work is related to COT-GAN in that we also use a distance metric as the discriminator which is the maximum mean discrepancy (MMD) calculated using the signature kernel proposed in [19]. However, as the computation of the signature kernel MMD does not require the use of a neural network, it belongs in the class of Score-based Generative Models (SGMs) described in [33] instead.
@@
The closest related work is [17] where the authors used a neural stochastic differential equation (SDE) as the base model which is a continuous time analogue of a recurrent neural network (RNN) with a stochastic evolution. One key difference to the approach in this paper is the distribution of the noise input to the generative model. The authors used independently distributed Gaussian noise for each time step while we use a moving average model to model the variance of the noise. This was found to be important in generating autocorrelation in the absolute returns of the generated sequences which is a stylised fact of financial time series.
@@
In terms of applying the generative model to downstream tasks, the authors of [20] used the GAN based generative model to create synthetic data, which are used for building or fine tune trading strategies. The strategies are based on data driven machine learning techniques such as neural networks and gradient boosting machines. We will do a similar portfolio management task using reinforcement learning trained on synthetic data created by the generative model
@@ 3.1 Reproducing Kernel Hilbert Space
The primary advantage of using a kernel is that there is a way to compute the result of the inner product without having to explicitly compute the feature map $\phi(x)$. This results in efficient computation despite the fact that the inner product space ${\mathcal{V}}$ may be of high dimension or even infinite dimensional. This is known as the kernel trick (see [35, Section 4]). Every kernel possesses a corresponding unique reproducing kernel Hilbert space (RKHS) (see [28]).
@@ 3.2 Kernel Two Sample Test
Firstly, by [28, Lemma 3.1], $\mu_{P},\mu_{Q}\in{\mathcal{H}}$. Moreover
@@
From [13, Lemma 6], let $X:=\{x_{1},\ldots,x_{m}\}$ and $Y:=\{y_{1},\ldots,y_{m}\}$ be samples from $P,Q\in{\mathcal{P}({\mathcal{X}})}$ respectively. Then an unbiased estimate of the $\text{MMD}^{2}$ is given by
@@ 3.3 Signature of a Path
The kernel two sample tests described in Section 3.2 are typically used for distributions over "static" i.e. non-causal values. However, we are interested in distributions over paths and to capture the time causalities of these path values. We can extend the test to distributions over paths by using the $\widehat{\text{MMD}^{2}}$ defined in equation (6) with $k$ being the signature kernel as demonstrated in [5]. Before we define the signature of a path which is a sequence of iterated integrals, we use an abbreviated notation for the integral. For a path ${\mathbf{x}}\in{\mathscr{P}}:=\{{\mathbf{x}}:[0,T]\to{\mathbb{R}}^{d}\}$ where at each time $t\in[0,T]$, ${\mathbf{x}}_{t}:=(x_{t}^{1},\ldots,x_{t}^{d})$ is the value of the path at time $t$, the integral of the path from time 00 to $t$ of asset $i\in\{1,\ldots,d\}$ is denoted as $S({\mathbf{x}})^{i}_{0,t}$.
@@
    Factorial decay of the signature terms [26, Proposition 2.2]: Assuming ${\mathbf{x}}$ is a path of finite total variation. For each $1\leq k\in{\mathbb{N}}$, the norm of the collection of terms of the signature of order $k$ (i.e. $k$ times iterated itegrals) is bounded as follows
@@
    Universal non-linearity [22, Theorem 2.1] :
    Let $D\subseteq{\mathscr{P}}$ be a compact subset [^3] Refer to [24, Section 8] on the topology of such a set of paths of the space of paths with bounded total variation from $[0,T]$ to ${\mathbb{R}}^{d}$ such that for any ${\mathbf{x}}\in D$, ${\mathbf{x}}_{0}=\mathbf{0}$ and at least one dimension of ${\mathbf{x}}$ is monotone (for uniqueness of the signature). Let $g:D\to{\mathbb{R}}$ be a continuous function. Then, for any $\epsilon>0$, there exists a linear functional $h$ such that for all ${\mathbf{x}}\in D$,
@@ 3.4 Signature Kernel
The rational quadratic kernel can be seen as a weighted sum of an infinite number of Gaussian kernels with different length scales where the squared inverse length scale follows a gamma distribution (see [38, Section 4.2]). The choice of the length scale parameter is important as it determines the smoothness of the kernel function. It is a hyperparameter that needs to be tuned for the specific problem at hand. The larger the length scale, the smoother the kernel function which means it is less sensitive to differences in the inputs. However, if the chosen length scale is too small then even the smallest differences arising from samples from the same distribution will be overemphasised.
The computation of the signature kernel from Definition 7 can be done efficiently using only evaluations of the static kernel on the paths ${\mathbf{x}}$ and ${\mathbf{y}}$ as shown in [19, Algorithm 3]  [^5] There is also a method to compute an untruncated signature kernel by numerically solving a PDE where the solution is equal to the untruncated signature kernel introduced in [30]. However, the MMD failed to converge in our experiments with this method.. The main hyperparameter is the truncation order $m$ of the signature. We modify the algorithm provided in https://github.com/tgcsaba/KSig for our specific use case.
@@ 4 Generative Model
The output of the generative model is the log return of the current time step, $r_{t_{i}}\in{\mathbb{R}}$. These returns are cumulatively summed to obtain the log price sequence. Once we have the generated log sequence, we perform the lead-lag and time augmentation on the sequence. With a batch of augmented sequences, we calculate the MMD between the generated sequences and the reference sequences from real data. We employ variants of stochastic gradient descent (see e.g. [11, Section 8.5]) to minimise the MMD between the generated sequences and the reference sequences from real data. Our objective is to have the generator output sequences that are close to the reference sequences, as measured by the MMD.
@@ 4.3 Noise
Before performing the inverse Lambert W transformation, we perform a preprocessing step for numerical reasons explained in [21]. First, we annualise the log returns by dividing them by the time elapsed in years. Then we normalise them to have zero mean and unit variance to get the annualised and normalised returns, $r_{Z}$. The inverse Lambert W transformation is then applied by estimating the parameters $\delta,\mu_{R}$ and $\sigma_{R}$ using iterative generalised method of moments [^9] Although $r_{Z}$ has been normalised with zero mean and unit varaiance. The estimation process will aim to produce a kurtosis of 3 for the transformed variables which will cause the value of $\mu_{R}$ and $\sigma_{R}$ to deviate from zero. from [10, Section 4.2]. Using the estimated parameters, we obtain the "Gaussianised" returns, $r_{W}=W(r_{Z})$.

### v3

@@ Abstract
Generating synthetic financial time series data that accurately reflects real-world market dynamics holds tremendous potential for various applications, including portfolio optimization, risk management, and large scale machine learning. We present an approach for training generative models for financial time series using the maximum mean discrepancy (MMD) with a signature kernel. Our method leverages the expressive power of the signature transform to capture the complex dependencies and temporal structures inherent in financial data. We employ a moving average model to model the variance of the noise input, enhancing the model’s ability to reproduce stylized facts such as volatility clustering. Through empirical experiments on S&P 500 index data, we demonstrate that our model effectively captures key characteristics of financial time series and outperforms a comparable GAN-based approach. In addition, we explore the application of the synthetic data generated to train a reinforcement learning agent for portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
@@ 1 Introduction
The availability of high-quality, realistic synthetic data for financial time series holds immense potential for various quantitative finance applications (see e.g. [3]). However, capturing the complex and nuanced characteristics of financial markets, such as volatility clustering, fat tails, and long-range dependencies, poses a significant challenge for traditional generative models. Recent advances in deep learning have introduced promising techniques, including Variational Autoencoders (VAEs) introduced in [30] and Generative Adversarial Networks (GANs) pioneered in [20], for generating synthetic financial data.
This paper proposes a generative model based on the maximum mean discrepancy (MMD) as defined in [21] with the use of signature kernels proposed by [31] to address these challenges. MMD provides a framework for comparing probability distributions, enabling effective training of the generative model. The signature kernel, derived from the signature transform, offers a powerful tool for capturing nuances of financial time series data. We further enhance the model’s ability to represent stylized facts by employing a moving average model to generate structured noise input, leading to more realistic synthetic data generation.
To demonstrate the effectiveness of our approach, we compare the characteristics of the generated sequence using our MMD approach against two alternative approaches. One approach from [59] uses causal optimal transport (COT) hence named COT-GAN, which is a type of Wasserstein distance that respects causality (see [4]). Another approach from [37], named Conditional Sig-WGAN, is similar to ours in that it also uses a MMD distance but does not first "lift" the sequence to a RKHS (see Section 3.4). We find that although both models are generally able to capture stylised facts of financial time series, there are some improvements in the statistical aspects of the MMD generated sequence. Furthermore, we explore the application of the generated synthetic data for training a reinforcement learning agent on the task of portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
The sections are organised as follows. Section 2 provides an overview of related works in the field of generative models for financial time series that use neural networks. Section 3 details the technical underpinnings behind the methodology. We describe our proposed generative model and the modelling of the noise variance in Section 4. Section 5 presents the experimental results, comparing our approach with the COT-GAN and Conditional Sig-WGAN, and evaluating the performance of the synthetic data generated for portfolio management using reinforcement learning. We also introduce a method to add robustness to the generative model in this section. Finally, Section 6 concludes the paper and discusses potential future research directions.
@@ 2 Related Works
Generative models are a type of statistical model that can be used to generate new data instances. They are especially useful in the financial domain, where they can be used to generate financial time series data. Financial time series data is a sequence of values that are measured at regular intervals, such as daily, weekly, or monthly. Generative models can be used to generate new time series data that is similar to the original data. This is useful for a variety of applications, including portfolio optimization, risk management, and privacy protection. In addition, generative models can be used to backtest trading strategies or serve as training data for machine learning models. Traditional statistical methods for modelling financial time series include autoregressive models, GARCH models, and stochastic volatility models. We focus instead on related works that also use neural networks as the backbone for the generative model.
@@
One of the earlier work on financial time series is from [32] which used restricted Boltzmann machines to generate foreign exchange rates. More recent advances in deep learning include approaches such as the Variational Autoencoders (VAEs) where the generative model is trained to maximise the evidence lower bound (ELBO) of the reference data as introduced in [30]. In [7], the authors trained VAEs to reproduce the log signatures of the reference data with a diagonal Gaussian distribution for the latent space. The choice of log signature is motivated by the fact that it is a unique feature transformation that captures the key characteristics of the path. The challenge lies in inverting the log signature to the original path, which is highly nontrivial (see e.g. [43]). In [56] the authors used autoencoders for dimensionality reduction of discrete local volatilities and utilises normalising flows to model the distribution of the latent variables while enforcing no-arbitrage constraints.
Another approach uses diffusion based models (see [24]). In [33], this was used to generate audio data in wave form. More recently, in [26], the FTS-Diffusion model was proposed to generate financial time series with irregular and scale-invariant patterns.
@@
An even more popular approach is the framework of Generative Adversarial Networks (GANs) pioneered in [20] to train neural networks generators of financial time series. One of the earliest works using this approach is [54] where the authors used multi-layer perceptrons and convolutional networks in the generator and discriminator. The QuantGAN from [55] focuses on the choice of neural network architecture. In particular, an architecture called Temporal Convolutional Networks (TCN) was used to model drift and volatility processes in order to generate samples. [17] further incorporates the attention mechanism with transformers into the network architecture under the GAN framework. One common issue with the GAN framework is that it is known to be exhibit instabilities during training. The authors of QuantGAN report that "training was very irregular and did not converge to a local optimum of the objective function", hence the need for model selection across multiple parameters saved during training.
@@
There are other GAN type approaches such as the TimeGAN from [60], which blends the GAN framework with autoencoders and supervised learning by simultaneously training the autoencoder to create the latent space and the generator and discriminator to model the dynamics of the latent space. TailGAN from [13] focuses on generating time series that preserve the tail risk features of the training data by exploiting the joint elicitability property of value-at-risk and expected shortfall. In [29], the authors use neural SDEs (see [28]), with the Wasserstein-1 distance to train the generator similar to the Wasserstein GAN from [2]. This falls under the GAN framework as the Wasserstein-1 distance is calculated using an optimisation based formulation that uses neural networks as function approximators. The COT-GAN from [59], uses causal Wasserstein or Causal Optimal Transport (COT) (see [4]) which adds an additional constraint to respect causality. Although this particular work was not applied to financial time series, it is an example of a GAN based approach that uses a distance metric as the discriminator exhibiting better stability during training. More recently, [1] proposed a time-causal VAE (TC-VAE) which uses VAEs and causal Wasserstein distance to train the generator for financial time series.
There are also works that use Wasserstein like distances as the scoring function for training the generator. In [45], the authors use the universality property of path signatures to replace the space of Lipschitz functions with linear functionals on the signature space in the Kantorovich-Rubinstein dual formulation of the Wasserstein distance. This switch is critical as it admits an analytical solution to the optimisation problem which means the training of the generator is no longer adversarial and belongs to the class of Score-based Generative Models (SGMs) described in [50]. The approach was further extended to the conditional setting with the Conditional Sig-WGAN in [37] and in [38]. In [5] the authors show that randomised signatures (see [11]), which is related to the signature transform and has the same expressive power while being finite-dimensional, also have the universality property. Therefore, it can be used to construct the Wasserstein distance as in [45] for training the generator.
@@
The closest related work is [27] where the authors also used neural SDEs as the base model and train the generator using a modified version of the MMD with the signature kernel. Similar to our approach, the authors also use the insight from [31] to first lift the path to the canonical feature space of a classical kernel such as the Gaussian kernel. One key difference to the approach in this paper is the distribution of the noise input to the generative model. It is typical in generative models to use independently distributed Gaussian noise for each time step. We use a moving average model to model the variance of the noise which is found to be important in generating autocorrelation in the absolute returns of the generated sequences which is a stylised fact of financial time series. The use of the moving average model is novel to the best of our knowledge.
@@
In terms of applying the generative model to downstream tasks, the authors of [10] simulate limit order book data while in [34], the authors create synthetic data for building or fine tune trading strategies. The strategies are based on data driven machine learning techniques such as neural networks and gradient boosting machines. We will do a similar portfolio management task using reinforcement learning trained on synthetic data created by the generative model. In particular, we look to generate synthetic data of longer time horizons which is another aspect that differentiates our approach from the existing works.
@@ 3.1 Reproducing Kernel Hilbert Space
The primary advantage of using a kernel is that there is a way to compute the result of the inner product without having to explicitly compute the feature map $\phi(x)$. This results in efficient computation despite the fact that the inner product space ${\mathcal{V}}$ may be of high dimension or even infinite dimensional. This is known as the kernel trick (see [52, Section 4]). Every kernel possesses a corresponding unique reproducing kernel Hilbert space (RKHS) (see [44]).
@@ 3.2 Kernel Two Sample Test
Firstly, by [44, Lemma 3.1], $\mu_{P},\mu_{Q}\in{\mathcal{H}}$. Moreover
@@
From [21, Lemma 6], let $X:=\{x_{1},\ldots,x_{m}\}$ and $Y:=\{y_{1},\ldots,y_{m}\}$ be samples from $P,Q\in{\mathcal{P}({\mathcal{X}})}$ respectively. Then an unbiased estimate of the $\text{MMD}^{2}$ is given by
@@ 3.3 Signature of a Path
The kernel two sample tests described in Section 3.2 are typically used for distributions over "static" i.e. non-causal values. However, we are interested in distributions over paths and to capture the time causalities of these path values. We can extend the test to distributions over paths by using the $\widehat{\text{MMD}^{2}}$ defined in equation (6) with $k$ being the signature kernel [9]. Before we define the signature of a path which is a sequence of iterated integrals, we use an abbreviated notation for the integral. For a path ${\mathbf{x}}\in{\mathscr{P}}:=\{{\mathbf{x}}:[0,T]\to{\mathbb{R}}^{d}\}$ where at each time $t\in[0,T]$, ${\mathbf{x}}_{t}:=(x_{t}^{1},\ldots,x_{t}^{d})$ is the value of the path at time $t$, the integral of the path from time 00 to $t$ of asset $i\in\{1,\ldots,d\}$ is denoted as $S({\mathbf{x}})^{i}_{0,t}$.
@@
    Factorial decay of the signature terms [42, Proposition 2.2]: Assuming ${\mathbf{x}}$ is a path of finite total variation. For each $1\leq k\in{\mathbb{N}}$, the norm of the collection of terms of the signature of order $k$ (i.e. $k$ times iterated itegrals) is bounded as follows
@@
    Universal non-linearity [36, Theorem 2.1] :
    Let $D\subseteq{\mathscr{P}}$ be a compact subset [^3] Refer to [40, Section 8] on the topology of such a set of paths of the space of paths with bounded total variation from $[0,T]$ to ${\mathbb{R}}^{d}$ such that for any ${\mathbf{x}}\in D$, ${\mathbf{x}}_{0}=\mathbf{0}$ and at least one dimension of ${\mathbf{x}}$ is monotone (for uniqueness of the signature). Let $g:D\to{\mathbb{R}}$ be a continuous function. Then, for any $\epsilon>0$, there exists a linear functional $h$ such that for all ${\mathbf{x}}\in D$,
@@ 3.4 Signature Kernel
The rational quadratic kernel can be seen as a weighted sum of an infinite number of Gaussian kernels with different length scales where the squared inverse length scale follows a gamma distribution (see [57, Section 4.2]). The choice of the length scale parameter is important as it determines the smoothness of the kernel function. It is a hyperparameter that needs to be tuned for the specific problem at hand. The larger the length scale, the smoother the kernel function which means it is less sensitive to differences in the inputs. However, if the chosen length scale is too small then even the smallest differences arising from samples from the same distribution will be overemphasised.
The computation of the signature kernel from Definition 7 can be done efficiently using only evaluations of the static kernel on the paths ${\mathbf{x}}$ and ${\mathbf{y}}$ as shown in [31, Algorithm 3]  [^5] There is also a method to compute an untruncated signature kernel by numerically solving a PDE where the solution is equal to the untruncated signature kernel introduced in [47]. However, the MMD failed to converge in our experiments with this method.. The main hyperparameter is the truncation order $m$ of the signature. We modify the algorithm provided in https://github.com/tgcsaba/KSig for our specific use case.
@@ 4 Generative Model
The output of the generative model is the log return of the current time step, $r_{t_{i}}\in{\mathbb{R}}$. These returns are cumulatively summed to obtain the log price sequence. Once we have the generated log sequence, we perform the lead-lag and time augmentation on the sequence. With a batch of augmented sequences, we calculate the MMD between the generated sequences and the reference sequences from real data. We employ variants of stochastic gradient descent (see e.g. [19, Section 8.5]) to minimize the MMD between the generated sequences and the reference sequences from real data. Our objective is to have the generator output sequences that are close to the reference sequences, as measured by the MMD.
@@ 4.2 The Cell State and Hidden State of the LSTM
Remark 1.
We emphasize that the conditioning described above is different from generating a conditional distribution such as in [37, Section 5] and [27, Section 3.4]. In both cases, the generator is trained to reproduce the conditional distribution of the future prices (or returns) given a specific past sequence. In [37], this is done by first assuming the true conditional expectation of the future price signature is a linear function of the past price signature and then using linear regression to estimate the parameters. Thereafter, the Monte Carlo estimate of the $l^{2}$ norm between the generator’s prediction given the past price serves as the loss function to train the generator. Using a linear estimator for the future path signature is no longer possible when we first lift the price path into a RKHS such as in our approach and in [27]. Therefore, the approach in [27] is to use a single historical sample of past and future sequence to estimate the score of multiple generated sequences conditioned on the same past sequence. In our approach, we estimate the score (MMD) by comparing a batch of generated sequences conditioned on a range of past sequences to the batch of future sequences resulting from the set of past sequences. We found that this approach combined with our choice of noise distribution described in Section 4.3 works well when generating long sequences (equal to about 1 calendar year of trading days) as we will show in Section 5.
@@ 4.3 Noise
Before performing the inverse Lambert W transformation, we perform a preprocessing step for numerical reasons explained in [35]. First, we annualise the log returns by dividing them by the time elapsed in years. Then we normalise them to have zero mean and unit variance to get the annualised and normalised returns, $r_{Z}$. The inverse Lambert W transformation is then applied by estimating the parameters $\delta,\mu_{R}$ and $\sigma_{R}$ using iterative generalised method of moments [^9] Although $r_{Z}$ has been normalised with zero mean and unit varaiance. The estimation process will aim to produce a kurtosis of 3 for the transformed variables which will cause the value of $\mu_{R}$ and $\sigma_{R}$ to deviate from zero. from [18, Section 4.2]. Using the estimated parameters, we obtain the "Gaussianised" returns, $r_{W}=W(r_{Z})$.